OpenStudy (anonymous):

[(sin(x))/(1-cos(x))]+[(sin(x))/(1+cos(x))]=2csc(x)

OpenStudy (anonymous):

These just keep getting better and better :D \[\Large \frac{\sin x}{1-\cos x}+\frac{\sin x}{1+\cos x}=2\csc x\]

OpenStudy (anonymous):

yes that is the equation. Please help me.

OpenStudy (anonymous):

. \[ \frac{\sin x}{1-\cos x}+\frac{\sin x}{1+\cos x}=\frac{(\sin x)(1+\cos x)}{(1-\cos x)(1+\cos x)}+\frac{(\sin x)(1-\cos x)}{(1-\cos x)(1+\cos x)}\]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Well, don't be a stranger ^.^ simplify it... Distribute what's to be distributed :D

OpenStudy (anonymous):

\[\frac{ \sin x+ \cos xsin x }{ (1-\cos x)(1+\cos x) } + \frac{-\sin x+ \cos xsinx }{(1-\cos x)(1+\cos x) }\]

OpenStudy (anonymous):

Wrong. The second fraction, why is the sin x negative? :/

OpenStudy (anonymous):

ohhhh so the second one looks like: \[\frac{ sinx -cosxsinx }{ (1-\cos x)(1+\cos x) }\]

OpenStudy (jdoe0001):

bbfoster123, keep in mind that (a-b)(1+b) = a^2-b^2 and that 1-cos^2 = sin^2

OpenStudy (anonymous):

So... \[\frac{ \sin x+ \cos x\sin x }{ (1-\cos x)(1+\cos x) } + \frac{\sin x- \cos x\sin x }{(1-\cos x)(1+\cos x) }\] They now have the same denominator, you can simply combine the numerators :) Do this now :D

OpenStudy (anonymous):

thanks @jdoe0001 @PeterPan the equation would now be: \[\frac{ \sin x + cosxsinx +\sin x - cosxsinx }{ (1-\cos x)(1+\cos x) }\]

OpenStudy (anonymous):

Yes... now, you can simplify the denominator as @jdoe0001 instructed (a - b)(a + b) = a² - b² And you can also simplify the numerator further~ by combining like terms ^.^ Do this now...

OpenStudy (anonymous):

so it is:\[\frac{ \sin^2x + \cos^2x \sin^2x }{ \sin^2 x }\]

OpenStudy (anonymous):

No... try again :) We're adding the numerators, not multiplying :P

OpenStudy (anonymous):

\[\frac{ 2\sin x }{ \sin^2x }\]

OpenStudy (anonymous):

Much better :) Can anything be cancelled out? ^.^

OpenStudy (anonymous):

sin x = 1/csc x

OpenStudy (anonymous):

Yeah, sure, but if you have \[\frac{2a}{a^2}\]doesn't one of the a's cancel out?

OpenStudy (anonymous):

yes they do. so it is sin x/ sin^2 x

OpenStudy (anonymous):

So... \[\huge \frac{2\sin x}{\sin^2 x}\]becomes...?

OpenStudy (anonymous):

\[\frac{ sinx }{ \sin^2 x }\]

OpenStudy (anonymous):

Why'd you remove the 2?

OpenStudy (anonymous):

arent we simplifying?

OpenStudy (anonymous):

Yes, we are. So how did the 2 disappear?

OpenStudy (anonymous):

\[\frac{ 2\sin x }{ \sin x }\]

OpenStudy (anonymous):

Now, you forgot that the sin at the bottom was squared.

OpenStudy (anonymous):

\[\frac{ 2\sin x }{ \sin ^2 x }\]

OpenStudy (anonymous):

Now... cancel out what CAN be cancelled out :)

OpenStudy (anonymous):

so its \[\frac{ \sin x }{ \sin x } = \frac{ 1/\csc }{ 1/\csc } \]

OpenStudy (anonymous):

then you mutiply and you get 2csc x

OpenStudy (anonymous):

If it works for you, then :)

OpenStudy (anonymous):

it doesn't work at all.

OpenStudy (anonymous):

hello...

OpenStudy (anonymous):

Because... as I said, you only have to cancel out. Don't convert it to csc yet.\[\frac{ 2\sin x }{ \sin ^2 x }\]

OpenStudy (anonymous):

the cancelling part confuses me.

OpenStudy (anonymous):

How would you simplify \[\large \frac{2a}{a^2}\]?

OpenStudy (anonymous):

\[\frac{ 2 }{ a }\]

OpenStudy (anonymous):

Well, I don't really see what's so hard about that, so how to simplify \[\Large \frac{ 2\sin x }{ \sin ^2 x }\]?

OpenStudy (anonymous):

\[\frac{ 2 }{ \sin x }\]

OpenStudy (anonymous):

See? Simple ^.^ Now what's 1/sin x?

OpenStudy (anonymous):

csc

OpenStudy (anonymous):

And there you have it :)

OpenStudy (anonymous):

Thank You!!!!!!